Dublin University Rifle Club

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Mark Dennehy 167 Applewood Hts Greystones Co. Wicklow Ireland

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Please find enclosed (finally! ☺) the prizes for the Trinity Top Ten for the year 2000-2001, as well as for the year 1999-2000. I’d like to personally apologise for these prizes not being made available sooner. Steps have been taken to ensure it will not happen in the next Top Ten, which I hope to see everyone enter again next year as it will be marking our 40^{th} anniversary.

Trinity Top Ten, 2000-2001

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10

Trinity Top Ten, 1999-2000 Gun Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Points Score  Place
Dave Cullen  RRPC  198  196  199  197  196  196  194  196  197  197  198  1
Ray Reilly  DURC  190  191  198  194  194  199  193  198  194  197  180  2
Ronan Power  RRPC  192  196  191  192  184  193  190  191  188  193  157  3
Agnes Graham  Comber  191  191  194  189  186  196  189  190  192  189  156  4
P.J. Harding  DRC  188  189  191  187  191  185  193  192  197  194  152  5
John Kennedy  FSC  187  191  193  193  188  189  192  193  188  192  151  6
J.K. Walsh  DRC  188  185  193  183  186  196  187  187  185  185  136  7

Note that according to the records of the Match Secretary for the 1999-2000 season, only seven people completed the Top Ten and were thus eligble

Mise le meas,

Mark Dennehy, DURC Secretary

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Dave Cooney Ardnaglass Corrig Road Dun Laoghaire

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Please find enclosed (finally! ☺) the prizes for the Trinity Top Ten for the year 2000-2001, as well as for the year 1999-2000. I’d like to personally apologise for these prizes not being made available sooner. Steps have been taken to ensure it will not happen in the next Top Ten, which I hope to see everyone enter again next year as it will be marking our 40^{th} anniversary.

Trinity Top Ten, 2000-2001

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10

Trinity Top Ten, 1999-2000 Gun Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Points Score  Place
Dave Cullen  RRPC  198  196  199  197  196  196  194  196  197  197  198  1
Ray Reilly  DURC  190  191  198  194  194  199  193  198  194  197  180  2
Ronan Power  RRPC  192  196  191  192  184  193  190  191  188  193  157  3
Agnes Graham  Comber  191  191  194  189  186  196  189  190  192  189  156  4
P.J. Harding  DRC  188  189  191  187  191  185  193  192  197  194  152  5
John Kennedy  FSC  187  191  193  193  188  189  192  193  188  192  151  6
J.K. Walsh  DRC  188  185  193  183  186  196  187  187  185  185  136  7

Mise le meas,

Mark Dennehy, DURC Secretary

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

John K. Walsh Carjon Lodge Cross Chapel Blessington Co. Wicklow

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Please find enclosed (finally! ☺) the prizes for the Trinity Top Ten for the year 2000-2001, as well as for the year 1999-2000. I’d like to personally apologise for these prizes not being made available sooner. Steps have been taken to ensure it will not happen in the next Top Ten, which I hope to see everyone enter again next year as it will be marking our 40^{th} anniversary.

Trinity Top Ten, 2000-2001

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10

Trinity Top Ten, 1999-2000 Gun Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Points Score  Place
Dave Cullen  RRPC  198  196  199  197  196  196  194  196  197  197  198  1
Ray Reilly  DURC  190  191  198  194  194  199  193  198  194  197  180  2
Ronan Power  RRPC  192  196  191  192  184  193  190  191  188  193  157  3
Agnes Graham  Comber  191  191  194  189  186  196  189  190  192  189  156  4
P.J. Harding  DRC  188  189  191  187  191  185  193  192  197  194  152  5
John Kennedy  FSC  187  191  193  193  188  189  192  193  188  192  151  6
J.K. Walsh  DRC  188  185  193  183  186  196  187  187  185  185  136  7

Mise le meas,

Mark Dennehy, DURC Secretary

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Agnes Graham 114 Towerview Crescent Bangor Co. Down BT19 6A2

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Please find enclosed (finally! ☺) the prizes for the Trinity Top Ten for the year 2000-2001, as well as for the year 1999-2000. I’d like to personally apologise for these prizes not being made available sooner. Steps have been taken to ensure it will not happen in the next Top Ten, which I hope to see everyone enter again next year as it will be marking our 40^{th} anniversary.

Trinity Top Ten, 2000-2001

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10

Trinity Top Ten, 1999-2000 Gun Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Points Score  Place
Dave Cullen  RRPC  198  196  199  197  196  196  194  196  197  197  198  1
Ray Reilly  DURC  190  191  198  194  194  199  193  198  194  197  180  2
Ronan Power  RRPC  192  196  191  192  184  193  190  191  188  193  157  3
Agnes Graham  Comber  191  191  194  189  186  196  189  190  192  189  156  4
P.J. Harding  DRC  188  189  191  187  191  185  193  192  197  194  152  5
John Kennedy  FSC  187  191  193  193  188  189  192  193  188  192  151  6
J.K. Walsh  DRC  188  185  193  183  186  196  187  187  185  185  136  7

Mise le meas,

Mark Dennehy, DURC Secretary

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Ivor Fletcher 127 Belfast Road Antrim Co. Antrim BT41 2AZ

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Please find enclosed (finally! ☺) the prizes for the Trinity Top Ten for the year 2000-2001, as well as for the year 1999-2000. I’d like to personally apologise for these prizes not being made available sooner. Steps have been taken to ensure it will not happen in the next Top Ten, which I hope to see everyone enter again next year as it will be marking our 40^{th} anniversary.

Trinity Top Ten, 2000-2001

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10

Trinity Top Ten, 1999-2000 Gun Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Points Score  Place
Dave Cullen  RRPC  198  196  199  197  196  196  194  196  197  197  198  1
Ray Reilly  DURC  190  191  198  194  194  199  193  198  194  197  180  2
Ronan Power  RRPC  192  196  191  192  184  193  190  191  188  193  157  3
Agnes Graham  Comber  191  191  194  189  186  196  189  190  192  189  156  4
P.J. Harding  DRC  188  189  191  187  191  185  193  192  197  194  152  5
John Kennedy  FSC  187  191  193  193  188  189  192  193  188  192  151  6
J.K. Walsh  DRC  188  185  193  183  186  196  187  187  185  185  136  7

Mise le meas,

Mark Dennehy, DURC Secretary

Dublin University Rifle Club ·Trinity College Dublin ·Dublin 2 ·Ireland

Karen Farrell 41 Deerpark Walk Blessington Co. Wicklow

Mark Dennehy, DURC Secretary

☎: +353-87-2258207

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten 2000,  5/11/2001

A Chara,

Please find enclosed (finally! ☺) the prizes for the Trinity Top Ten for the year 2000-2001, as well as for the year 1999-2000. I’d like to personally apologise for these prizes not being made available sooner. Steps have been taken to ensure it will not happen in the next Top Ten, which I hope to see everyone enter again next year as it will be marking our 40^{th} anniversary.

Trinity Top Ten, 2000-2001

Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Score  Place
Andrew McFarland  EARC  24  24  20  25  25  19  26  23  19  26  231  1
Gary Duff  Comber  26  26  26  24  21  25  21  26  12  12  219  2
John Walsh  DRC  19  18  16  26  26  23  24  24  20  22  218  3
Robert Clarke  EARC  23  25  22  17  19  20  23  19  23  18  209  4
Richard Farrell  RRPC  20  14  25  23  16  21  22  22  21  25  209  4
Agnes Graham  Comber  15  16  17  21  20  26  18  25  26  24  208  6
Tom McKnight  EARC  17  23  18  18  13  15  20  13  25  13  175  7
Trevor McFarland  EARC  21  19  9  20  15  12  16  15  24  20  171  8
Ronan Power  RRPC  11  13  12  13  24  16  25  17  16  16  163  9
Elizabeth Rodgers  EARC  12  8  10  14  10  22  15  21  22  23  157  10

Trinity Top Ten, 1999-2000 Gun Scores

Name  Club  Rnd 1  Rnd 2  Rnd 3  Rnd 4  Rnd 5  Rnd 6  Rnd 7  Rnd 8  Rnd 9  Rnd 10  Final Points Score  Place
Dave Cullen  RRPC  198  196  199  197  196  196  194  196  197  197  198  1
Ray Reilly  DURC  190  191  198  194  194  199  193  198  194  197  180  2
Ronan Power  RRPC  192  196  191  192  184  193  190  191  188  193  157  3
Agnes Graham  Comber  191  191  194  189  186  196  189  190  192  189  156  4
P.J. Harding  DRC  188  189  191  187  191  185  193  192  197  194  152  5
John Kennedy  FSC  187  191  193  193  188  189  192  193  188  192  151  6
J.K. Walsh  DRC  188  185  193  183  186  196  187  187  185  185  136  7

Mise le meas,

Mark Dennehy, DURC Secretary